# KAGNNs: Kolmogorov-Arnold Networks meet Graph Learning

### v3

@@ Results.
Table 5 illustrates the average classification accuracies and the corresponding standard deviations of the three models on the different datasets. We observe that GIN and both KAGIN  models achieve similar levels of performance on all datasets except ENZYMES. In particular, RBF-KAGIN  and BS-KAGIN  both have the best accuracy on $2$ datasets each, and GIN on $1$ dataset. Regarding the GCN variants, we observe that the KAN-based variants now have a significant advantage over the linear GCN, confirming that a single KAN layer is more expressive than a simple fully-connexted layer. BS-KAGCN  and RBF-KAGCN  obtain the best overall accuracy on $1$ dataset each, while the vanilla GCN performs often significantly poorly.
These experiments show that, on a graph-level task, which requires significantly more power than node-level tasks, KAN-based models seem to perform better than their MLP-based counterparts. While this difference is not significant among GIN-like architectures, it appears that KAN-based GCNs benefit from the higher expressivity in a single layer, and attain performances much closer to their GIN-like counterparts, sometimes even outperforming them.
@@ 4.3 Graph Regression
We experiment with two molecular datasets: (1) ZINC-12K [42], and (2) QM9 [43]. ZINC-12K consists of $12,000$ molecules. The task is to predict the constrained solubility of molecules, an important property for designing generative GNNs for molecules. The dataset is already split into training, validation and test sets ($10,000$, $1,000$ and $1,000$ graphs in the training, validation and test sets, respectively). QM9 contains approximately $134,000$ organic molecules. Each molecule consists of Hydrogen, Carbon, Oxygen, Nitrogen, and Flourine atoms and contain up to 9 heavy (non Hydrogen) atoms. The task is to predict $12$ target properties for each molecule. The dataset was divided into a training, a validation, and a test set according to a $80\%/10\%/10\%$ split.
@@
We perform grid search to select values for the different hyperparameters. We choose the number of hidden layers from $\{2,3,4\}$ for GIN and $\{1,2,3\}$ for KAN. We use different ranges for each model was because we wanted to limit the size advantage of KAN against MLP, since a layer with the same width of KAN has more parameters than an MLP one). For both models we search a learning rate from $\{10^{-3},10^{-4}\}$. For GIN, we choose the hidden dimension size from $\{32,64,128,256,512,1024\}$, while for KAGIN, we choose it from $\{4,8,16,32,64,128,256\}$. For KAGIN, we also select the grid size from $\{1,3,5,8,10\}$ and for BS-KAGIN the spline order from $\{3,5\}$. To produce graph representations, we use the sum operator. The emerging graph representations are finally fed to a 2-layer MLP (for GIN) or a 2-layer KAN (for KAGIN) which produces the output. We set the batch size equal to $128$ for all models. We train each model for $1,000$ epochs by minimizing the mean absolute error (MAE). We use the Adam optimizer for model training [39]. We also use early stopping with a patience of $20$ epochs. For ZINC-12K, we also use an embedding layer that maps node features into $100$-dimensional vectors. We choose the configuration that achieves the lowest validation error. Once the best configuration is found, we run $10$ experiments and report the average performance on the test set. For both datasets and models, we set the number of message-passing layers to $4$. On QM9, we performed a joint regression of the $12$ targets.
@@ 5 Analysis of Size and Time
5.1 Selected models
Following the above results, and it is legitimate to wonder whether the performance of the KAN models comes from a larger size than their MLP-based counterparts. We show in this section that it is not the case. Using the graph-classification models trained for the results provided in 4.2, we display here three relations: the size-to-performance relation, the size-to-training time relation and the training time-to-performance relation.
Figures 2 and 4 show these relations for datasets NCI1 and IMDB-MULTI respectively. We also show in 8 and 8 the numerical values associated to these graphs. The rest of the graph classification datasets are provided in App. 5.2. A first thing that we can see from both figures is that, for the same parameter count, B-splines-based KANs are significantly slower than both other architectures. RBF-based KANs are much faster, showing that this implementation succeeds in alleviating the computational bottleneck induced by the sequential computation necessary for B-splines. Finally, MLPs are the fastest for equal parameter counts. While these results were expected, we notice in both the tables and figures that RBF-KAGINmodels end up significantly smaller than their MLP-based counterparts, while maintaining equal or better performance, and about equal training time. On the other hand, the B-Splines based method is significantly slower, by sometimes a 5 fold factor. This establishes the RBF-based architecture as the best compromise between scalability and performance, although further work could be made in studying the performance of very small KAN-based against MLP-based models.
@@ 5.1 Selected models
Table 7: Time and parameter sizes for selected models, NCI1
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.794 $\pm$ 0.016 572338.4 2.220
RBF-KAGIN 0.783 $\pm$ 0.017 139168.0 0.358
GIN 0.783 $\pm$ 0.024 1337369.8 0.359
BS-KAGCN 0.716 $\pm$ 0.031 90196.5 0.752
RBF-KAGCN 0.640 $\pm$ 0.075 166960.5 0.356
GCN 0.604 $\pm$ 0.046 379634.0 0.277
Table 8: Time and parameter sizes for selected models, IMDB-MULTI
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.499 $\pm$ 0.033 296418.8 0.367
RBF-KAGIN 0.488 $\pm$ 0.043 48705.1 0.169
GCN 0.503 $\pm$ 0.038 330987.6 0.146
BS-KAGIN 0.500 $\pm$ 0.042 46052.3 0.216
RBF-KAGIN 0.485 $\pm$ 0.038 40118.5 0.161
MLP 0.462 $\pm$ 0.046 111911.4 0.143
5.2 Theoretical models
We give in Table 9 the training time per epoch for different configurations of the BS-KAGIN, RBF-KAGIN and GIN models on a graph-level task. We observe that for a given number of message-passing layers and hidden dimension size, GIN is the fastest, followed by RBF-KAGIN. BS-KAGIN is noticeably slower than the others. It suffers doubly: its parameter count increases much quicker than both other models for same hyperparameters; but even for similar numbers of parameters, BS-KAGIN is by far the slowest. Spline order seems to have a larger impact on running time than grid size. This shows that RBFs do alleviate some of the computational bottlenecks of the B-splines from the original KAN paper. Table 16 in Appendix C shows similar observations on a node-level task.
@@ 5.2 Theoretical models
6 Discussion
Prominent application domains.
KANs, both in our work and in the recent literature, seem to be a promising alternative to MLPs for regression tasks. They rely on combinations of continuous functions which could explain their improved performance for continuous outputs. Moreover, this property and their empirical performance in regression tasks renders them an ideal candidate for physics-inspired architectures [44], which address problems of high importance, from climate forecasting [45] to atomic simulations [46] and rely heavily on graph-based architectures.
Explainability matters.
The potential of explainability for KAN-based models is arguably better compared to that of MLPs with post-hoc methods (e.g. Shapley values, gradient-based), since closed-form symbolic formulas can be drawn from the former, providing complete overview of the underlying computations. Hence, KAN-based graph models could potentially prove more friendly to practitioners from application fields, such as biology and physics, who tend to be skeptical of the black-box approach.
GNNs for fixed graphs inspired from KANs.
Each KAN architecture consists of a set of neurons. Each one of those neurons is connected to other neurons. Thus, KANs can be naturally modeled as graphs where nodes represent neurons and edges represent the connections between them. Each layer of a KAN is actually a complete bipartite graph, while a KAN has learnable activation functions on the edges of its associated graph. There actually exist datasets that contain multiple graphs which share the same graph structure, but differ in the node and edge features. Two examples of such datasets are the MNIST and CIFAR10 graph classification datasets [47]. Since the graph structure is fixed (and identical for all graphs), we could associate each edge of the graph with one or more learnable activation functions (shared across graphs), similar to what KANs do. This would allow the model to learn different functions for each edge of the graph, leading potentially to high-quality node and graph representations.
Advantages and future work.
Finally, we discuss potential advantages that KANs might have over MLPs, and leave their investigation for future work. First, their ability to accurately fit smooth functions could prove relevant on datasets where variables interact with some regular patterns. Second, their interpretability could be leveraged to provide explanations on learned models, giving insights into the nature of interactions among entities. Finally, a thorough study of the effect of the different hyperparameters could be pursued, allowing to fully exploit the splines or radial basis functions while retaining small networks.
7 Conclusion
In this paper, we have investigated the potential of Kolmogorov-Arnold networks in graph learning tasks. Since the KAN architecture is a natural alternative to the MLP, we developed two GNN architectures, KAGCN and KAGIN, respectively analogous to the GCN and GIN models. We then compared those architectures against each other in both node- and graph-level tasks. For graph and node classification, there does not appear to be a clear winner. For graph regression, our results indicate that KAN has an advantage over MLP, with the B-spline variant proving superior in terms of performance. Nonetheless, this gain comes at the cost of computational efficiency. In this regard, the RBF-based vairant offers a slightly lesser performance improvement, but proves much more efficient in terms of training time.
Appendix A Further Graph Classification Experiments
Besides the graph classification experiments that are presented in the main paper, we also experimented with the Peptides-func dataset, a graph classification dataset from the Long Range Graph Benchmark [48]. The dataset consists of $15,535$ graphs where the average number of nodes per graph is equal to $150.94$. We evaluated GIN, BS-KAGIN and RBF-KAGIN on this dataset. We performed a grid search where for all models we chose the learning rate from $\{0.001,0.0001\}$, the number of hidden layers of MLPs and KANs from $\{1,2,3\}$ and the hidden dimension from $\{64,128\}$. We set the number of neighborhood aggregation layers to $5$ for all models. For both BS-KAGIN and RBF-KAGIN, grid size was chosen from $\{3,5\}$, and for BS-KAGIN, spline order was chosen from $\{3,5\}$. Each experiment was run $4$ times with $4$ different seeds and Table 10 illustrates the average precision achieved by the three models.
Table 10: Average Precision ($\pm$ standard deviation) of the KAGIN and GIN models on the Peptides-func dataset.
Peptides-func
GIN 55.01 $\pm$ 0.75
BS-KAGIN
58.71
$\pm$ 1.02
RBF-KAGIN 57.68 $\pm$ 2.05
Appendix B Time/Performance/Accuracy analysis
Figures 6 to 14 and tables 11 to 15 regarding the time/size/accuracy analysis on the graph classification task.
Table 11: Time and parameter sizes for selected models, IMDB-BINARY
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.734 $\pm$ 0.038 148236.3 0.225
RBF-KAGIN 0.720 $\pm$ 0.036 168383.6 0.129
GIN 0.731 $\pm$ 0.038 1172520.5 0.110
BS-KAGCN 0.708 $\pm$ 0.037 39053.5 0.150
RBF-KAGCN 0.726 $\pm$ 0.044 37322.5 0.109
GCN 0.687 $\pm$ 0.034 118826.3 0.099
Table 12: Time and parameter sizes for selected models, ENZYMES
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.427 $\pm$ 0.085 450136.6 0.336
RBF-KAGIN 0.412 $\pm$ 0.079 115346.8 0.061
GIN 0.533 $\pm$ 0.117 1623033.2 0.048
BS-KAGCN 0.577 $\pm$ 0.076 83636.3 0.103
RBF-KAGCN 0.594 $\pm$ 0.053 125854.4 0.050
GCN 0.221 $\pm$ 0.053 280040.0 0.039
Table 13: Time and parameter sizes for selected models, DD
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.754 $\pm$ 0.032 169417.4 1.206
RBF-KAGIN 0.757 $\pm$ 0.029 147005.1 0.236
GIN 0.748 $\pm$ 0.044 168957.4 0.144
BS-KAGCN 0.733 $\pm$ 0.042 110278.6 0.596
RBF-KAGCN 0.683 $\pm$ 0.043 110899.2 0.185
GCN 0.632 $\pm$ 0.036 342023.2 0.225
Table 14: Time and parameter sizes for selected models, MUTAG
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.851 $\pm$ 0.058 236235.8 0.056
RBF-KAGIN 0.855 $\pm$ 0.051 62530.0 0.020
GIN 0.853 $\pm$ 0.069 848213.0 0.014
BS-KAGCN 0.755 $\pm$ 0.077 19496.1 0.022
RBF-KAGCN 0.735 $\pm$ 0.110 47769.4 0.017
GCN 0.563 $\pm$ 0.143 72072.2 0.012
Table 15: Time and parameter sizes for selected models, PROTEINS_full
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.741 $\pm$ 0.034 321367.7 0.472
RBF-KAGIN 0.733 $\pm$ 0.034 51498.8 0.096
MLP 0.710 $\pm$ 0.027 942781.0 0.073
BS-KAGCN 0.756 $\pm$ 0.034 37688.7 0.133
RBF-KAGCN 0.742 $\pm$ 0.035 48950.4 0.078
GCN 0.605 $\pm$ 0.020 8199.8 0.061
Figure 9: Parameters count/size/performance comparison, GIN, DD
Figure 10: Parameters count/size/performance comparison, GCN, DD
Appendix C Training Times - Node Classification

### v4

@@ Results.
Table 4 illustrates the average classification accuracies and the corresponding standard deviations of the three models on the different datasets. We observe that B-splines variants win on $4$ datasets, and the MLP-based GIN leads on $3$, albeit by small margins wrt. the standard deviation. We notice that, while the MLP-based GCN performs poorly (predictably so, since graph-level tasks often require higher complexity than node-classification ones), the KAN-based GCN perform significantly better. This tends to confirm that a single KAN layer has much more representation power than a fully-connected layer.
@@ 5.4 Graph Regression
We experiment with two molecular datasets: (1) ZINC-12K (Irwin and Shoichet, 2005), and (2) QM9 (Ramakrishnan et al., 2014). ZINC-12K consists of $12,000$ molecules. The task is to predict the constrained solubility of molecules, an important property for designing generative GNNs for molecules. The dataset is already split into training, validation and test sets ($10,000$, $1,000$ and $1,000$ graphs in the training, validation and test sets, respectively). QM9 contains approximately $134,000$ organic molecules. Each molecule consists of Hydrogen, Carbon, Oxygen, Nitrogen, and Flourine atoms and contain up to 9 heavy (non Hydrogen) atoms. The task is to predict $12$ target properties for each molecule. The dataset was divided into a training, a validation, and a test set according to a $80\%/10\%/10\%$ split.
@@
We perform grid search to select values for the different hyperparameters. We choose the number of hidden layers from $\{2,3,4\}$ for GIN and $\{1,2,3\}$ for KAN. We use different ranges for each model was because we wanted to limit the size advantage of KAN against MLP, since a layer with the same width of KAN has more parameters than an MLP one). For both models we search a learning rate from $\{10^{-3},10^{-4}\}$. For GIN, we choose the hidden dimension size from $\{32,64,128,256,512,1024\}$, while for KAGIN, we choose it from $\{4,8,16,32,64,128,256\}$. For KAGIN, we also select the grid size from $\{1,3,5,8,10\}$ and for BS-KAGIN the spline order from $\{3,5\}$. To produce graph representations, we use the sum operator. The emerging graph representations are finally fed to a 2-layer MLP (for GIN) or a 2-layer KAN (for KAGIN) which produces the output. We set the batch size equal to $128$ for all models. We train each model for $1,000$ epochs by minimizing the mean absolute error (MAE). We use the Adam optimizer for model training (Kingma and Ba, 2015). We also use early stopping with a patience of $20$ epochs. For ZINC-12K, we also use an embedding layer that maps node features into $100$-dimensional vectors. We choose the configuration that achieves the lowest validation error. Once the best configuration is found, we run $10$ experiments and report the average performance on the test set. For both datasets and models, we set the number of message-passing layers to $4$. On QM9, we performed a joint regression of the $12$ targets.
@@ 6 Analysis of Size and Time
6.1 Setting
Following the above results, and it is legitimate to wonder whether the performance of the KAN models comes from a larger size than their MLP-based counterparts. We show in this section that it is not the case. Using the graph-classification models trained for the results provided in 5.3, we display here three relations: the size-to-performance relation, the size-to-training time relation and the training time-to-performance relation. Note that these models are the ones selected by the hyperparameter search algorithm, thus exhibiting the best validation performance throughout the architecture-optimization process.
We thus notice an interesting tradeoff, where RBF-based KANs tend to train models with good performance, but smaller size than their MLP-based counterparts, allowing them to train almost as quickly. Meanwhile BS-based KANs tend to be slower.
6.2 Discussion
Figures 1 shows these relations on dataset NCI1 for the GIN-based models, a setting where the three models have a similar performance. We also show in table 6 the associated numerical values for this dataset. The rest of the graph classification datasets and models are provided in App. C.1. We notice that, for the same number of parameters, MLPs are the fastest, followed closely by RBF-based KANs; then BS-based KANs are significantly slower than both others. This confirms the fact that the RBF-based representation alleviates greatly the computational bottlenecks of B-splines. Nonetheless, we also notice that, despite the hyperparameter search setting allowing them a large size, the performance-driven model selection converged towards smaller sizes for KAN-based models than for MLPs. This allows the RBF-based version of KAGNN models to train in time similar to their MLP-based counterpart. Nonetheless, the BS-based variant remains significantly slower in almost all settings.
@@ 6.2 Discussion
Table 6: Average time and parameter sizes for selected models, NCI1
Model Parameter count Accuracy time per epoch (s)
GCN 316588.40 0.60 $\pm$ 0.04 0.31
BS-KAGCN 182464.80 0.74 $\pm$ 0.04 0.94
RBF-KAGCN 117043.10 0.73 $\pm$ 0.03 0.37
GAT 1566334.00 0.60 $\pm$ 0.08 0.47
BS-KAGAT 1109714.80 0.76 $\pm$ 0.04 1.42
RBF-KAGAT 1468335.80 0.73 $\pm$ 0.04 0.53
GIN 1557454.00 0.79 $\pm$ 0.02 0.39
BS-KAGIN 532684.50 0.80 $\pm$ 0.01 1.70
RBF-KAGIN 107075.40 0.79 $\pm$ 0.02 0.37
7 Discussion and Perspectives
Prominent application domains.
KANs, both in our work and in the recent literature, seem to be a promising alternative to MLPs for regression tasks. They rely on combinations of continuous functions which could explain their improved performance for continuous outputs. Moreover, this property and their empirical performance in regression tasks renders them an ideal candidate for physics-inspired architectures (Karniadakis et al., 2021), which address problems of high importance, from climate forecasting (Kashinath et al., 2021) to atomic simulations (Batzner et al., 2022) and rely heavily on graph-based architectures.
Explainability matters.
Recent works on KANs argue that their potential for explainability is higher than that of MLPs with post-hoc methods (e.g. Shapley values, gradient-based) (Liu et al., 2025), since closed-form symbolic formulas can be drawn from the former (Liu et al., 2024), providing complete overview of the underlying computations. Hence, KAN-based models could prove more friendly to practitioners from application fields, such as biology and physics, who tend to be skeptical of the black-box approach.
GNNs for fixed graphs inspired from KANs.
Each KAN architecture consists of a set of neurons. Each one of those neurons is connected to other neurons. Thus, KANs can be naturally modeled as graphs where nodes represent neurons and edges represent the connections between them. Each layer of a KAN is actually a complete bipartite graph, while a KAN has learnable activation functions on the edges of its associated graph. There actually exist datasets that contain multiple graphs which share the same graph structure, but differ in the node and edge features. Two examples of such datasets are the MNIST and CIFAR10 graph classification datasets (Dwivedi et al., 2023). Since the graph structure is fixed (and identical for all graphs), we could associate each edge of the graph with one or more learnable activation functions (shared across graphs), similar to what KANs do. This would allow the model to learn different functions for each edge of the graph, leading potentially to high-quality node and graph representations.
Advantages and future work.
Finally, we discuss potential advantages that KANs might have over MLPs, and leave their investigation for future work. First, their ability to accurately fit smooth functions could prove relevant on datasets where variables interact with some regular patterns. Second, their interpretability could be leveraged to provide explanations on learned models, giving insights into the nature of interactions among entities. Another study direction could be to extend the KAN architecture to more GNN models (e.g. GAT, GraphSAGE, Graph Transformer) to develop a wider KAN-based arsenal for graph tasks. Finally, a thorough study of the effect of the different hyperparameters could be pursued, allowing to fully exploit the splines or radial basis functions while retaining small networks.
8 Conclusion
In this paper, we have investigated the potential of Kolmogorov-Arnold networks in graph learning tasks. Since the KAN architecture is a natural alternative to the MLP, we developed three GNN architectures, KAGCN, KAGAT and KAGIN, respectively analogous to the GCN, GAT and GIN models. We then compared those architectures against each other in both node- and graph-level tasks. For link prediction, as well as graph and node classification, there does not appear to be a clear winner. For graph regression, our results indicate that KAN has an advantage over MLP, with the B-spline variant proving superior in terms of performance. Nonetheless, this gain comes at the cost of computational efficiency. In this regard, the RBF-based variant offers a lesser performance improvement, but proves much more efficient in terms of training time.
Acknowledgements
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation and partner Swedish universities and industry. Part of the experiments was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. Part of the experiments was also enabled by the Berzelius resource, which was made possible through application support provided by National Supercomputer Centre at Linköping University.
Appendix A Further Graph Classification Experiments
Besides the graph classification experiments that are presented in the main paper, we also experimented with the Peptides-func dataset, a graph classification dataset from the Long Range Graph Benchmark [Dwivedi et al., 2022]. The dataset consists of $15,535$ graphs where the average number of nodes per graph is equal to $150.94$. We evaluated GIN, BS-KAGIN and RBF-KAGIN on this dataset. We performed a grid search where for all models we chose the learning rate from $\{0.001,0.0001\}$, the number of hidden layers of MLPs and KANs from $\{1,2,3\}$ and the hidden dimension from $\{64,128\}$. We set the number of neighborhood aggregation layers to $5$ for all models. For both BS-KAGIN and RBF-KAGIN, grid size was chosen from $\{3,5\}$, and for BS-KAGIN, spline order was chosen from $\{3,5\}$. Each experiment was run $4$ times with $4$ different seeds and Table 7 illustrates the average precision achieved by the three models.
Table 7: Average Precision ($\pm$ standard deviation) of the KAGIN and GIN models on Peptides-func
Peptides-func
GIN 55.01 $\pm$ 0.75
BS-KAGIN
58.71
$\pm$ 1.02
RBF-KAGIN 57.68 $\pm$ 2.05
Appendix B Hyperparameter search ranges
Table 8: Hyperparameter ranges for node classification (datasets Cora, Cornell, Texas, Wisconsin)
Model Learning Rate Hidden Layers Hidden Dim. Grid Size Spl. Order Dropout
GIN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[1,256\right]$ NA NA $\left[0,0.9\right]$
BS-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[1,128\right]$ $\left[1,8\right]$ $\left[1,3\right]$ $\left[0,0.9\right]$
RBF-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,128\right]$ $\left[2,32\right]$ NA $\left[0,0.9\right]$
Table 9: Hyperparameter ranges for node classification (datasets CiteSeer, Ogbn-Arxiv, Actor)
Model Learning Rate Hidden Layers Hidden Dim. Grid Size Spl. Order Dropout
GIN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[1,64\right]$ NA NA $\left[0,0.9\right]$
BS-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,32\right]$ $\left[1,8\right]$ $\left[1,3\right]$ $\left[0,0.9\right]$
RBF-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,32\right]$ $\left[2,4\right]$ NA $\left[0,0.9\right]$
Table 10: Hyperparameter ranges for link prediction
Model Learning Rate Hidden Layers Hidden Dim. Grid Size Spl. Order
GIN $\{1e-2,1e-3,1e-4\}$$\{2,4\}$$\{16,512\}$ NA NA
[…]
RBF-KAGIN/KAGCN $\{1e-2,1e-3\}$$\{2,4\}$$\{16,512\}$$\{2,16\}$ NA
Table 11: Hyperparameter ranges for graph classification
Appendix C Time/Performance/Accuracy analysis
Figures 2 to 22 and tables 12 to 17 regarding the time/size/accuracy analysis on the graph classification task.
Table 12: Average time and parameter sizes for selected models, MUTAG
Model Parameter count Accuracy time per epoch (s)
GCN 66421.40 0.60 $\pm$ 0.09 0.02
BS-KAGCN 31991.40 0.72 $\pm$ 0.07 0.03
RBF-KAGCN 67221.20 0.78 $\pm$ 0.09 0.02
GAT 1162504.80 0.58 $\pm$ 0.12 0.02
BS-KAGAT 521626.00 0.83 $\pm$ 0.05 0.04
RBF-KAGAT 1119860.90 0.76 $\pm$ 0.11 0.02
GIN 840858.00 0.86 $\pm$ 0.06 0.02
BS-KAGIN 145455.00 0.86 $\pm$ 0.08 0.05
RBF-KAGIN 82645.40 0.85 $\pm$ 0.05 0.02
Table 13: Average time and parameter sizes for selected models, DD
Model Parameter count Accuracy time per epoch (s)
GCN 274320.80 0.66 $\pm$ 0.06 0.24
BS-KAGCN 110847.70 0.73 $\pm$ 0.04 0.93
RBF-KAGCN 110976.90 0.67 $\pm$ 0.06 0.24
GAT 224323.60 0.72 $\pm$ 0.06 0.22
BS-KAGAT 1375974.80 0.78 $\pm$ 0.02 2.36
RBF-KAGAT 514070.80 0.73 $\pm$ 0.05 0.38
GIN 281992.60 0.75 $\pm$ 0.04 0.15
BS-KAGIN 189140.10 0.76 $\pm$ 0.05 2.24
RBF-KAGIN 62553.50 0.76 $\pm$ 0.03 0.24
Table 14: Average time and parameter sizes for selected models, PROTEINS_full
Model Parameter count Accuracy time per epoch (s)
GCN 14055.70 0.63 $\pm$ 0.04 0.09
BS-KAGCN 26241.00 0.76 $\pm$ 0.03 0.15
RBF-KAGCN 77950.00 0.74 $\pm$ 0.04 0.11
GAT 1217234.00 0.60 $\pm$ 0.01 0.14
BS-KAGAT 377421.60 0.73 $\pm$ 0.02 0.28
RBF-KAGAT 406871.00 0.72 $\pm$ 0.04 0.13
GIN 935245.00 0.71 $\pm$ 0.03 0.10
BS-KAGIN 206523.00 0.73 $\pm$ 0.03 0.36
RBF-KAGIN 76445.40 0.74 $\pm$ 0.03 0.14
Table 15: Average time and parameter sizes for selected models, ENZYMES
Model Parameter count Accuracy time per epoch (s)
GCN 230071.20 0.21 $\pm$ 0.04 0.06
BS-KAGCN 102457.20 0.59 $\pm$ 0.04 0.12
RBF-KAGCN 105379.80 0.57 $\pm$ 0.08 0.08
GAT 2193344.40 0.18 $\pm$ 0.01 0.10
BS-KAGAT 734566.80 0.51 $\pm$ 0.12 0.23
RBF-KAGAT 292237.60 0.50 $\pm$ 0.07 0.09
GIN 778435.40 0.55 $\pm$ 0.06 0.07
BS-KAGIN 199432.90 0.44 $\pm$ 0.08 0.21
RBF-KAGIN 135783.50 0.45 $\pm$ 0.09 0.10
Table 16: Average time and parameter sizes for selected models, IMDB-BINARY
Model Parameter count Accuracy time per epoch (s)
GCN 106247.80 0.68 $\pm$ 0.04 0.15
BS-KAGCN 34710.10 0.73 $\pm$ 0.05 0.19
RBF-KAGCN 40357.50 0.73 $\pm$ 0.04 0.17
GAT 1482555.60 0.67 $\pm$ 0.04 0.18
BS-KAGAT 391572.00 0.72 $\pm$ 0.04 0.25
RBF-KAGAT 308826.00 0.71 $\pm$ 0.05 0.18
GIN 769820.20 0.73 $\pm$ 0.03 0.16
BS-KAGIN 144101.00 0.73 $\pm$ 0.04 0.30
RBF-KAGIN 159608.70 0.70 $\pm$ 0.04 0.20
Table 17: Average time and parameter sizes for selected models, IMDB-MULTI
Model Parameter count Accuracy time per epoch (s)
GCN 83223.60 0.49 $\pm$ 0.04 0.23
BS-KAGCN 50801.40 0.50 $\pm$ 0.04 0.29
RBF-KAGCN 25513.00 0.50 $\pm$ 0.04 0.25
GAT 838081.80 0.48 $\pm$ 0.05 0.25
BS-KAGAT 129868.00 0.50 $\pm$ 0.04 0.33
RBF-KAGAT 340916.00 0.48 $\pm$ 0.04 0.27
GIN 875383.20 0.50 $\pm$ 0.04 0.24
BS-KAGIN 195207.70 0.50 $\pm$ 0.04 0.38
RBF-KAGIN 40460.40 0.49 $\pm$ 0.04 0.26
Figure 3: Parameters count/size/performance comparison, GAT, MUTAG
Figure 5: Parameters count/size/performance comparison, GCN, DD
Figure 6: Parameters count/size/performance comparison, GAT, DD
Figure 7: Parameters count/size/performance comparison, GIN, DD
@@ Appendix C Time/Performance/Accuracy analysis
Figure 9: Parameters count/size/performance comparison, GAT, NCI1
Figure 10: Parameters count/size/performance comparison, GIN, NCI1
Figure 12: Parameters count/size/performance comparison, GAT, PROTEINS_full
Figure 15: Parameters count/size/performance comparison, GAT, ENZYMES
Figure 18: Parameters count/size/performance comparison, GAT, IMDB-BINARY
Figure 21: Parameters count/size/performance comparison, GAT, IMDB-MULTI
@@
C.1 Training Times - Graph Classification
We give in Table 18 the training time per epoch for different configurations of the BS-KAGIN, RBF-KAGIN and GIN models on a graph-level task. We observe that for a given number of message-passing layers and hidden dimension size, GIN is the fastest, followed by RBF-KAGIN. BS-KAGIN is noticeably slower than the others. It suffers doubly: its parameter count increases much quicker than both other models for same hyperparameters; but even for similar numbers of parameters, BS-KAGIN is by far the slowest. Spline order seems to have a larger impact on running time than grid size. This shows that RBFs do alleviate some of the computational bottlenecks of the B-splines from the original KAN paper. Table LABEL:tab:timesNodeClassif in Appendix C.2 shows similar observations on a node-level task.
@@ C.1 Training Times - Graph Classification
C.2 Training Times - Node Classification